OpenStudy (anonymous):

Mary and Janet are running in a race. Mary runs at 7 mph. Janet runs at 5mph. If they start at the same time, how long will it be before they're 1.5 miles apart?

OpenStudy (mertsj):

7t+5t=1.5

OpenStudy (anonymous):

The answer says 45 minutes. How do you get that from 7t+5t=1.5

OpenStudy (mertsj):

What did you get when you solve that equation?

OpenStudy (anonymous):

t=.125

OpenStudy (anonymous):

OH is that hours?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

and you convert it into minutes?

OpenStudy (anonymous):

duh! Thank you!!

OpenStudy (mertsj):

Which is 7.5 minutes. Not 45

OpenStudy (anonymous):

damn. yeah

OpenStudy (anonymous):

maybe the answer is wrong let me recheck

OpenStudy (anonymous):

it says 45 minutes in the book ugh whatever

OpenStudy (mertsj):

Did you post the problem correctly? Maybe there is a typo.

OpenStudy (mertsj):

If they are running in the same direction then we should write 7t-5t=1.5 2t=1.5 t=.75 hours = 45 min